OpenStudy (anonymous):

how the integration by parts formula related to the product rule of differentiation?

OpenStudy (lgbasallote):

Product rule: \[(uv)' = uv' + vu'\] in other words \[\frac{d}{dx} (uv) = udv + vdu\] subtract vdu from both sides \[\implies \frac{d}{dx} (uv) - vdu = udv\] integrate both sides \[\implies \int \frac{d}{dx} (uv) - \int vdu = \int udv\] simplify \[\implies uv - \int vdu = \int udv\] look familiar?